OpenStudy (anonymous):

Find the derivative function of cos^3[sin(2x)] Find the derivative function of cos^3[sin(2x)] @Mathematics

OpenStudy (anonymous):

\[-6 \sin (\sin (2 x)) \cos (2 x) \cos ^2(\sin (2 x)) \]